OpenStudy (anonymous):

75% of a number is 120. What is 40% of the number?

OpenStudy (anonymous):

id like to know how to get the answer to! (^.^)

OpenStudy (anonymous):

\[\frac{120}{x}=\frac{75}{100}\] Solve for x to find the number. Now, solve for y.\[\frac{y}{x}=\frac{40}{100}\]

OpenStudy (anonymous):

thank you! (^.^)

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

Or solve one at a time 0.75 N = 120 N = 120/.75 = 160 0.40 N = (0.40)(160) = 64 answer